OpenStudy (melissa_412):

Can someone please help me with this quick 8th grade math question? Fiona and Iliana went to a going-out-of-business sale at a local video store. The store was advertizing all HD videos on sale for one price and all classic videos for a different price. Fiona bought 5 HD videos and 2 classic videos for $31. Iliana bought 5 classic videos and 3 HD videos for $30. Which system of equations can be used to find the prices of the classic and HD videos?

OpenStudy (melissa_412):

Hello

OpenStudy (melissa_412):

Hi

OpenStudy (xguardians):

5h + 2c = $31 3h + 5c = $30 Choose an equation and solve for a variable 3h + 5c = 30 5c = 30 - 3h c = (30-3h)/5 Substitute into another equation 5h + 2(30-3h)/5 = 31 5h + (60-6h)/5 = 31 5h + 12 - 6h/5 = 31 5h - 6h/5 = 19 3.8h = 19 h = 5 Resubstitute into original equation 5(5) + 2c = 31 25 + 2c = 31 2c = 6 c = 3 H = HD videos = $5.00 each C = classic videos = $3.00 each

OpenStudy (melissa_412):

I don't get it @xGuardians ?

OpenStudy (xguardians):

What don't you understand?

OpenStudy (melissa_412):

I mean the question is asking me for a system of equations.

OpenStudy (melissa_412):

Oh never mind

OpenStudy (xguardians):

5h + 2c = $31 3h + 5c = $30 Is that all you needed?

OpenStudy (melissa_412):

Can you help with one more? Thuy wrote the system of equations. 24x+36y=72 -16x+8y=80 If the second equation is multiplied by 3, what should the first equation be multiplied by to eliminate the x variable by addition? @xGuardians

OpenStudy (melissa_412):

Yea lol

OpenStudy (xguardians):

So all you really care about in this equation is the x values. 3(-16x) = -48x 24x * what value = 48x So -48x and 48x cancel out.

OpenStudy (melissa_412):

So 0?

OpenStudy (xguardians):

Nope.

OpenStudy (xguardians):

24 * ? = 48

OpenStudy (melissa_412):

)=

OpenStudy (melissa_412):

2

OpenStudy (xguardians):

Yeah. So the answer is two.

OpenStudy (melissa_412):

Sorry I'm just innately bad at math

OpenStudy (xguardians):

I'll write it out. NP.

OpenStudy (melissa_412):

Which operations would create an equivalent system of equations with opposite like terms? 3x-3y=3 4x+5y=13 @xGuardians

OpenStudy (melissa_412):

=)

OpenStudy (xguardians):

2(24x+36y=72) 3(-16x+8y=80) BREAK 48x + 72y = 144 -48x + 24y = 240 ______________________ 0x + 96y = 384 So when you add them together you get 0x

OpenStudy (melissa_412):

Oh sorry I forgot to include the answer choices A. The first equation can be multiplied by 5 and the second equation by 3. B. The first equation can be multiplied by –4 and the second equation by –3. C. The first equation can be multiplied by 4 and the second equation by –4. D. The first equation can be multiplied by 3 and the second equation by 5.

OpenStudy (melissa_412):

?

OpenStudy (xguardians):

A I think, when you apply it it will make the y-values cancel out.

OpenStudy (melissa_412):

Ok thanks so much that was the last one